OpenStudy (anonymous):

find the present value PV of the given investment. round your answer to the nearest cent. An investment earns 5% per year and is worth $1,000 after 6 months.

OpenStudy (anonymous):

is it 1001.3?

OpenStudy (anonymous):

because\[1000/1+0.05*6\]

OpenStudy (wolf1728):

Total = princ * (1 + rate)^n where 'n' is years Total = 1,000 * (1.05)^.5 Total = 1,000 * 1.0246950766 Total = 1,024.70

OpenStudy (wolf1728):

If it was for 1 full year then Total = 1,000 * 1.05^1 Total = 1,050.00 SO, roughly after 6 months it earns $24 interest

OpenStudy (anonymous):

Isn't it \[1000/(1.05)*6\]

OpenStudy (anonymous):

for the 6 months

OpenStudy (wolf1728):

That equals 158.73 Unless you mean (1,000/1.05)*6 which equals 5,714.29

OpenStudy (anonymous):